A company produces car engines with the following technology: Q=5KL where K denotes the number of conveyor belts and L denotes the number of working teams. The rent for one conveyor belt is 10000 per week and one team of workers costs 5000 per week. Additionally, each engine requires materials costing 2000. Imagine that the number of conveyor belts is fixed to five in the short-run Which equation denotes the short-run total cost of producing Q engines? Wählen Sie eine Antwort: O a. TC(Q) 2200+50000/Q b. TC(Q) 2000Q+100000 c. TC(Q) 2000+100000/Q d. TC(Q) 2200Q+50000
Answer #1

Rent = r = 10000

Wages = w = 5000

K = 5

Material cost = 2000 ( Variable cost )

Production function = Q = 5KL

Q = 5*5*L

L = Q/25

Total cost = Variable cost + fixed cost

Variable cost = Material cost * Production function+ wL

VC = 2000Q + (5000*Q/25)

VC = 2200Q

Fixed cost = FC = rK

FC = 10000*5 = 50000

TC = Total cost = VC + FC

TC = 2200Q+50000

Answer: d